OpenStudy (anonymous):

I have a question if anyone can help me. Thanks! The width of a rectangle is 1 less than twice its length. If the area of the rectangle is 180 cm^2, what is the length of the diagonal?

OpenStudy (ranga):

Assume Length is L Width = 2L - 1 Area = LW = L(2L - 1) = 180 Solve for L. Find W. Then the diagonal will be sqrt(L^2 + W^2)